OpenStudy (anonymous):

Is the function convergent or divergent? Prove your work! \sum_{0}^{\infty}(1+\sin(n))/(10^n) Award given at the end! :D

OpenStudy (anonymous):

\[\sum_{0}^{\infty}(1+\sin(n))/(10^n) \]

OpenStudy (freckles):

hint: \[\sin(n) \le 1\]